OpenStudy (anonymous):

(Sq rt 4x+3) = -2 please help show all work

OpenStudy (anonymous):

start by raising both sides to the 2nd power this will clear the radical and then you solve for x like normal

OpenStudy (anonymous):

need all steps please, did that and got wrong answer.

OpenStudy (dumbcow):

No solution the sqrt of a number must be positive

OpenStudy (anonymous):

answer should be "null"

OpenStudy (anonymous):

sq -2= positive 4?! where am i going wrong?

OpenStudy (anonymous):

raise the -2 to the 2nd power and you get 4

OpenStudy (anonymous):

i have the following: 4x+3=4 , 4x= 1, x= 1/3

OpenStudy (anonymous):

now the radical is removed from the other side because you did that also imaginarly

OpenStudy (anonymous):

1/4 th, sorry.. not 1/3

OpenStudy (dumbcow):

that is correct so x = 1/4 could be a solution however sqrt(4) does not equal -2

OpenStudy (anonymous):

so, if I get an answer.. how is it a null? like the book answer says?

OpenStudy (dumbcow):

by plugging it back into original equation and check if equality holds if it doesn't then its no solution or "null"

OpenStudy (anonymous):

ohh, right.. i forgot that ty , ty